In Macroland, the GDP deflator for 2014 is 105 with 2013 as the base year. Real GDP in 2014 equalled $210 billion. Therefore, nominal GDP in 2014 equalledwhile nominal G DP in 2013 equalled
Answer #1

Answer to blank 1: $220.5 billion

Answer to blank 2: $200 billion

Explanation:

Nominal GDP in 2014 = (105 * 210) / 100 = $220.5 billion

Nominal gdG in 2013 = (210 / 105) * 100 = $200 billion